# HC3 Plus: A Semantic-Invariant Human ChatGPT Comparison Corpus

### v2

@@ Abstract
ChatGPT has gained significant interest due to its impressive performance, but people are increasingly concerned about its potential risks, particularly around the detection of AI-generated content (AIGC), which is often difficult for untrained humans to identify. Current datasets utilized for detecting ChatGPT-generated text primarily center around question-answering, yet they tend to disregard tasks that possess semantic-invariant properties, such as summarization, translation, and paraphrasing. Our primary studies demonstrate that detecting model-generated text on semantic-invariant tasks is more difficult. To fill this gap, we introduce a more extensive and comprehensive dataset that considers more types of tasks than previous work, including semantic-invariant tasks. In addition, the model after a large number of task instruction fine-tuning shows a strong powerful performance. Owing to its previous success, we further instruct fine-tuning Tk-instruct and built a more powerful detection system.
@@ 1 Introduction
In recent years, artificial intelligence has made great progress. Since the release of ChatGPT, large-scale language models (LLMs) fine-tuned on GPT3.5 with Reinforcement Learning from Human Feedback (RLHF) have received extensive attention Christiano et al. (2017); Stiennon et al. (2020). ChatGPT has shown strong performance on various tasks and is good at generating human-like text. Guo et al. (2023) show by the Turing test that it is difficult for amateurs (who have never heard of ChatGPT) to identify which texts are generated by ChatGPT or humans.
As more and more AI-generated content appears on social media, this could increase the risk of AI-generated disinformation and harmful text spreading due to the difficulty for humans to distinguish them Guo et al. (2023). Therefore, how to detect the text generated by the large language model has become a recent research hotspot. In fact, a good detector can be used for information supervision and accountability, making the source of information clearer. Guo et al. (2023) proposed HC3 a dataset containing text generated by ChatGPT and human-written text for training a detector. They first collected nearly 40k questions and responses from human experts via social media and wiki text. Then use ChatGPT to generate answers to these questions. Next, using these collected data, they established two detection methods, including training a logistic regression model based on GLTR Gehrmann et al. (2019) to complete the classification and using RoBERTa Liu et al. (2019) to train a binary classification model.
@@
While their detector excels in performance on their test set, it is primarily focused on the question-answering (QA) task. In the QA task, the responses produced by the language model merely need to answer the requested question, without strict adherence to the semantics of the raw question. However, for tasks such as summarization, translation, and paraphrasing, the LLM must carefully consider the semantic nuances of the input sentences. The generated sentences must stay true to the meaning of the raw sentences without deviating from them. In other words, the output sentences need to maintain semantic invariance with the raw sentences. At the same time, the raw sentence used as input can be used as a reference for the output of the LLM. In the example shown in Figure 1, when comparing summarization with QA, there is an amount of overlapping vocabulary between the source sentence and the target sentence, and they have similar semantics. This may make it more difficult to recognize that the text is generated by the LLM.
@@
We focus on translation, summarization, and paraphrasing datasets to confirm the above perceptions. We first select several commonly used human-annotated datasets, then take raw sentences as input of ChatGPT, and collect the generated output. In order to verify whether the detector can distinguish these data well, we mix human-labeled target sentences in the datasets and the sentences generated by ChatGPT as the input of the detector, and let it determine which texts were generated by ChatGPT or human. We use chatgpt-detector-RoBERTa-chinese111https://huggingface.co/Hello-SimpleAI/chatgpt-detector-roberta-chinese to detect Chinese and chatgpt-detector-RoBERTa222https://huggingface.co/Hello-SimpleAI/chatgpt-qa-detector-roberta-chinese to detect English respectively, which were proposed by Guo et al. (2023). As shown in Table 1, we find that current detectors have difficulty distinguishing the source of sentences. The prediction results are mainly presented in two cases. In the first case, as shown in the translation dataset WMT’19 enzh results, the detector performs poorly in terms of precision and recall for both classes. The second case, as shown in the summarization dataset CNN/DailyMail See et al. (2017) results, where the detector is always biased towards predicting text as written by a human, exhibits very high recall in the human class but very low recall in the ChatGPT class. For direct comparison with detection results on the QA samples of HC3, we propose HC3-Paraphrase, which contains paraphrases of HC3 questions generated using ChatGPT. Although ChatGPT takes roughly the same input (both question text) for QA and paraphrasing tasks, as shown in Table 1, the detection performance on the HC3-Paraphrase is greatly degraded compared to the HC3. This result validates our hypothesis that current detectors struggle to detect semantic-invariant tasks.
@@
In order to fill the gap of HC3 under semantic-invariant tasks, we extend HC3 and propose a larger ChatGPT-generated text dataset covering translation, summarization, and paraphrasing tasks, called HC3 Plus. Furthermore, language models for fine-tuning instructions on a large number of tasks have shown impressive capabilities and generalization. We combine the data of HC3 and the data we proposed and further perform instruction fine-tuning in the Tk-instruct Wang et al. (2022) model to train a detector. Experimental results show that the model trained by our instruction fine-tuning method has a stronger detection ability than the model proposed by Guo et al. (2023), and achieves a substantial improvement in performance on semantic-invariant tasks.
Our main contribution can be summarized as follows:
@@
We propose a more powerful detector based on instruction fine-tuning.
@@ 2 Dataset Construction
The LCSTS Hu et al. (2015) is a large-scale Chinese short text summarization dataset utilizing naturally annotated web resources on Sina Weibo, which is a Chinese social media like Twitter.
@@
The WMT (Workshop on Machine Translation) translation dataset is a widely used collection of data for machine translation. In this paper, we follow HC3 and mainly focus on the construction of English and Chinese detectors. Therefore, we consider the translation set to include WMT’19 EnZh, WMT’19 ZhEn, WMT’16 RoEn, WMT’16 Deen and WMT’14 FrEn.
@@
We call GPT-3.5-Turbo via the OpenAI API to generate target texts. We call our proposed dataset HC3-SI (HC3 semantic-invariance), which is roughly twice the size of HC3. In Appendix B, we detail the composition and size of the dataset. In addition, we divide 10% from the training data of HC3 as its validation set. Then, we combined the data of HC3-SI and HC3, which we called HC3 Plus for convenience.
@@ 3 Method
The language model fine-tuned on a large number of instructions shows strong generalization ability on unseen tasks. We propose instructDGGC, to utilize extensive instruction-finetuned LLM for Detecting ChatGPT-Generated Content. Our method is based on the Tk-instruct model and further instruction fine-tuning on the HC3 Plus dataset. InstructDGGC is a generative model for discriminating whether an input text is generated by ChatGPT given in-context instructions (task definition and few-shot).
@@
POSITIVE EXAMPLES are input samples and their correct output, denoted as . In this paper, we give 2-shot each time, one is a sample generated by ChatGPT, and the other is a sample written by a human.
@@ 4 Experiment
We followed the experimental setup of Guo et al. (2023) and use HC3 and HC3 Plus to train Roberta-based detectors, resulting in detectors called Roberta-HC3 and Roberta-HC3 Plus, respectively. Subsequently, we train the InstructDGGC model, which is divided into two steps. First, we follow the Tk-instruct to conduct instruction tuning on a large number of tasks, and then further fine-tune the model on HC3 Plus. In this section, we introduce the experimental details.
@@ 4.1 Experiment Setting
Before training the detector, we need to perform instruction fine-tuning on a large number of tasks, so that our generative model has good instruction-following ability. For the English model, we use allenai/Tk-instruct-base-def-pos333https://huggingface.co/allenai/tk-instruct-base-def-pos released by Wang et al. (2022), which has been fine-tuned on 757 English tasks. However, they did not publish a base-size model that can handle Chinese. Therefore, we conduct instruction tuning a language model LM using instruction-equipped data. We used the code released by Wang et al. (2022), followed their settings, and trained an instruction fine-tuning model on Super-NaturalInstructions(including 24 Chinese tasks) based on mt5-base Xue et al. (2021).
We then further conduct instruction fine-tuning on the HC3 Plus data. At this stage, for both Chinese and English models, we use the same parameter configuration and we train for 4 epochs with a learning rate of 1e-4 and a batch size of 32.
@@ 4.2 Experiment Result and Analysis
First of all, we focus on the RoBERTa-based model. The RoBERTa-HC3 performed poorly on the data we proposed, which shows that the model cannot handle the semantic-invariant data generated by ChatGPT. We then use HC3 Plus to train RoBERTa. For semantic-invariant data, we found that in addition to the decline in the Chinese translation data, the model performance of other data has been significantly improved. We observe that for the translation task, ChatGPT generates sentences that are very similar to the target sentences in the dataset, compared to summarization and paraphrasing. Therefore, it is more difficult for the model to detect translation data, even if a lot of training data is used. In Appendix C we give examples and a more detailed analysis. Furthermore, we observed that RoBERTa-HC3 trained only on HC3 data has a stronger detection ability on HC3 than RoBERTaa-HC3 Plus. This shows that training a detector for a single task is easier than training a detector that needs to consider various task data.
Then, we pay attention to InstructDGGC. As shown in Table 2, compared with RoBERTa-HC3 Plus, we find a significant improvement in English data compared to RoBERTa-HC3 Plus, which demonstrates good generalization of language models fine-tuned on a large number of instructions. However, on the Chinese data, except for the obvious improvement in the translation data, the results of the other parts are basically the same. We think that the reason for the limited improvement is that the instruction-following ability of the model is insufficient due to a lack of data for fine-tuning Chinese instructions. At the same time, compared with HC3-Plus, InstructDGGC performs better on translation data, which also proves that our model is more robust.
Finally, we focus on the overall score. Compared to RoBERTa-HC3 Plus, InstructDGGC achieved better overall performance, with an improvement of 1.8% for English data and 0.58% for Chinese data. This demonstrates the effectiveness of our method.
@@ 5 Conclusion
In this work, we experimentally demonstrate that detecting ChatGPT-generated text becomes more difficult on semantic-invariant tasks, for which we propose a more extensive and comprehensive dataset HC3 Plus. And we use the instruction fine-tuned model to achieve a detector and experimental results show that our detector outperforms the RoBERTa-based method proposed by previous work. We hope that our analysis of semantic-invariant text detection can provide some insights for future work.
@@ 6 Limitations
Although our experiments find that detecting semantic invariant text is more challenging for detectors, and our dataset only considered the current version of ChatGPT, it is still in continuous iterative improvement. This may result in our dataset lagging behind the constantly evolving versions of ChatGPT. We will further investigate the impact of the iteration of ChatGPT versions in our future work.
@@ Appendix B HC3-SI Dataset Description
As shown in Table 3, we give the number of samples in the data set HC3-SI. For Chinese data, the number of samples in the train/val/test set is 42708/4746/22516. For English data, the number of samples in the train/val/test set The sample sizes are 95745/10641/38142 respectively. The total sample size of our proposed HC3-SI is 210000, nearly twice that of HC3. In addition, our data does not overlap with HC3, and they can be used together.

### v4

@@ Abstract
ChatGPT has garnered significant interest due to its impressive performance; however, there is growing concern about its potential risks, particularly in the detection of AI-generated content (AIGC), which is often challenging for untrained individuals to identify. Current datasets used for detecting ChatGPT-generated text primarily focus on question-answering tasks, often overlooking tasks with semantic-invariant properties, such as summarization, translation, and paraphrasing. In this paper, we demonstrate that detecting model-generated text in semantic-invariant tasks is more challenging. To address this gap, we introduce a more extensive and comprehensive dataset that incorporates a wider range of tasks than previous work, including those with semantic-invariant properties. In addition, instruction fine-tuning has demonstrated superior performance across various tasks. In this paper, we explore the use of instruction fine-tuning models for detecting text generated by ChatGPT.
@@ 1 Introduction
In recent years, artificial intelligence has made significant advancements. Since the release of ChatGPT, large-scale language models (LLMs) fine-tuned on GPT-3.5 using Reinforcement Learning from Human Feedback (RLHF) have garnered extensive attention. ChatGPT has demonstrated strong performance across various tasks and excels at generating human-like text. As shown in Guo et al. (2023), the Turing test reveals that it is challenging for individuals unfamiliar with ChatGPT to distinguish between texts generated by ChatGPT and those written by humans.
As more and more AI-generated content appears on social media, it could increase the risk of AI-generated disinformation and harmful text spreading due to the difficulty for humans to distinguish them Guo et al. (2023). Therefore, how to detect the text generated by the large language model has become a recent research hotspot. In fact, a good detector can be used for information supervision and accountability, making the source of information clearer. Guo et al. (2023) introduced HC3, a dataset containing both ChatGPT-generated text and human-written text, designed for training AI text detection models. They first collected nearly questions and responses from human experts through social media and wiki text. ChatGPT was then used to generate answers to these questions. Using this collected data, they established two detection methods: training a logistic regression model based on GLTR Gehrmann et al. (2019) for classification and using RoBERTa Liu et al. (2019) to train a binary classification model.
While their detector excels in performance on their test set, it is primarily focused on the question-answering (QA) task. In the QA task, the responses produced by the language model merely need to answer the requested question, without strict adherence to the semantics of the raw question. However, for tasks such as summarization, translation, and paraphrasing, the LLM must carefully consider the semantic nuances of the input sentences. The generated sentences must stay true to the meaning of the raw sentences without deviating from them. In other words, the output sentences need to maintain semantic invariance with the raw sentences. At the same time, the raw sentence used as input can be used as a reference for the output of the LLM. In the example shown in Figure 1, when comparing summarization with QA, there is an amount of overlapping vocabulary between the source sentence and the target sentence, and they have similar semantics. That may make it more difficult to recognize that the text is generated by the LLM.
@@
We focus on translation, summarization, and paraphrasing datasets to validate the aforementioned observations. We begin by selecting several commonly used human-annotated datasets and inputting their raw sentences into ChatGPT to generate corresponding outputs. To verify whether the detector can effectively distinguish between these types of data, we mix the human-labeled target sentences from the datasets with the sentences generated by ChatGPT. These combined sentences are then provided as input to the detector, which determines whether each text is generated by ChatGPT or written by a human. We use chatgpt-detector-RoBERTa-chinese111https://huggingface.co/Hello-SimpleAI/chatgpt-detector-roberta-chinese to detect Chinese and chatgpt-detector-RoBERTa222https://huggingface.co/Hello-SimpleAI/chatgpt-qa-detector-roberta-chinese to detect English respectively, as proposed by Guo et al. (2023). As shown in Table 1, we find that current detectors have difficulty distinguishing the source of sentences. The prediction results are mainly presented in two cases. In the first case, as shown in the translation dataset WMT’19 enzh results, the detector performs poorly in terms of precision and recall for both classes. The second case, as shown in the summarization dataset CNN/DailyMail See et al. (2017) results, where the detector is always biased towards predicting text as written by a human, exhibits very high recall in the human class but very low recall in the ChatGPT class. For direct comparison with detection results on the QA samples of HC3, we propose HC3-Paraphrase, which contains paraphrases of HC3 questions generated using ChatGPT. Although ChatGPT takes roughly the same input (both question text) for QA and paraphrasing tasks, as shown in Table 1, the detection performance on the HC3-Paraphrase is greatly degraded compared to the HC3. The result validates our hypothesis that current detectors struggle to detect semantic-invariant tasks.
@@
In order to fill the gap of HC3 under semantic-invariant tasks, we extend HC3 and propose a larger ChatGPT-generated text dataset covering translation, summarization, and paraphrasing tasks, called HC3 Plus. In addition, language models fine-tuned on a large number of tasks have demonstrated impressive abilities and generalization Wang et al. (2022). In this paper, to explore the effectiveness of instruction fine-tuning models in detecting machine-generated text, we investigate the use of the Tk-instruct Wang et al. (2022) model for detecting text generated by ChatGPT. Our main contribution can be summarized as follows:
@@
We use instruction fine-tuning to train a stronger ChatGPT-based text generation detection model.
@@ 2 Dataset Construction
The LCSTS Hu et al. (2015) is a large-scale Chinese short-text summarization dataset utilizing naturally annotated web resources on Sina Weibo, which is a Chinese social media like Twitter.
@@
The WMT translation dataset is a widely used collection of data for machine translation. In this paper, we follow HC3 and mainly focus on the construction of English and Chinese detectors. Therefore, we consider the translation set to include WMT’19 EnZh, WMT’19 ZhEn, WMT’16 RoEn, WMT’16 Deen and WMT’14 FrEn.
@@
We use GPT-3.5-Turbo-0301 via the OpenAI API to generate target texts. Our proposed dataset, HC3-SI (HC3 Semantic-Invariance), is approximately twice the size of HC3. Details about the composition and size of HC3-SI can be found in Appendix B. We then combine HC3-SI with HC3 to create a dataset referred to as HC3 Plus for convenience.
@@ 3 Method
The language model fine-tuned on a large number of instructions shows strong generalization ability on unseen tasks. We propose instructDGGC, to utilize extensive instruction fine-tuning LLM for Detecting ChatGPT-Generated Content. Our method is based on the Tk-instruct model and further instruction fine-tuning on the HC3 Plus dataset. InstructDGGC is a generative model for discriminating whether an input text is generated by ChatGPT given in-context instructions (task definition and few-shot).
@@
POSITIVE EXAMPLES are input samples and their correct output, denoted as . In this paper, we give -shot each time, one is a sample generated by ChatGPT, and the other is a sample written by a human.
@@ 4 Experiment
First, we follow Guo et al. (2023) to use the Roberta-Base333https://huggingface.co/FacebookAI/roberta-base model for training the detector as a baseline. We train two RoBERTa models based on HC3 and HC3 Plus, named RoBERTa-HC3 and RoBERTa-HC3 Plus, respectively. Subsequently, we train the InstructDGGC model, which is divided into two steps. First, we follow the Tk-instruct to conduct instruction tuning on a large number of tasks, and then further fine-tune the model on HC3 Plus. In this section, we introduce the experimental details.
@@ 4.1 Experiment Setting
Before training the detector, we need to perform instruction fine-tuning on a large number of tasks, so that our generative model has good instruction-following ability. For the English model, we use Tk-instruct-base-def-pos444https://huggingface.co/allenai/tk-instruct-base-def-pos released by Wang et al. (2022), which has been fine-tuned on 757 English tasks. However, they did not publish a base-size model that can handle Chinese. Therefore, we conduct instruction tuning a language model LM using instruction-equipped data. We used the code released by Wang et al. (2022), followed their settings, and trained an instruction fine-tuning model on Super-NaturalInstructions(including 24 Chinese tasks) based on mt5-base Xue et al. (2021).
We then further conduct instruction fine-tuning on the HC3 Plus data. At this stage, for both Chinese and English models, we use the same parameter configuration and we train for epochs with a learning rate of and a batch size of .
@@ 4.2 Experiment Result and Analysis
First, we focus on the RoBERTa-based model. RoBERTa-HC3 performed poorly on our proposed dataset, indicating that the model struggles to handle semantic-invariant data generated by ChatGPT. We then used HC3 Plus to train RoBERTa. For semantic-invariant data, we observed that, aside from a decline in performance on the Chinese translation data, the model’s performance on other datasets improved significantly. We noted that for the translation task, ChatGPT generates sentences that are much more similar to the target sentences in the dataset compared to tasks like summarization and paraphrasing. As a result, detecting AI-generated content in translation data is more challenging for the model, even with a substantial amount of training data. Appendix C provides examples and a more detailed analysis. Additionally, we observed that RoBERTa-HC3, when trained only on HC3 data, demonstrates stronger detection capabilities on HC3 compared to RoBERTa-HC3 Plus. This suggests that training a detector for a specific task is easier than training one that needs to accommodate various types of task data.
Next, we focus on InstructDGGC. As shown in Table 2, we observe a significant improvement in performance on English data compared to RoBERTa-HC3 Plus, demonstrating the strong generalization capabilities of language models fine-tuned on a large number of instructions. However, for Chinese data, aside from a noticeable improvement in translation tasks, the results for other tasks remain largely unchanged. We believe the limited improvement is due to the model’s insufficient instruction-following capability, likely stemming from a lack of fine-tuning data for Chinese instructions. Moreover, compared to HC3 Plus, InstructDGGC performs better on translation data, further confirming that it is more robust.
Finally, we focus on the overall score. Compared to RoBERTa-HC3 Plus, InstructDGGC achieved better overall performance, with an improvement of % for English data and % for Chinese data. This indicates that fine-tuning instruction models for detecting generated text is a promising method.
@@ 5 Conclusion
In this work, we experimentally demonstrate that detecting ChatGPT-generated text becomes more difficult on semantic-invariant tasks, for which we propose a more extensive and comprehensive dataset HC3 Plus. And we use the instruction fine-tuning model to achieve a detector and experimental results show that our detector outperforms the RoBERTa-based method proposed by previous work. We hope that our analysis of semantic-invariant text detection can provide some insights for future work.
@@ 6 Limitations
While our experiments show that detecting semantic-invariant text is more challenging for detectors, our dataset only considers the current version of ChatGPT, i.e., GPT-3.5-Turbo-0301. As ChatGPT undergoes continuous iterative improvement, our dataset may become outdated compared to the evolving versions of ChatGPT. In future work, we plan to further investigate the impact of these iterations on detection performance.
@@ Appendix B HC3-SI Dataset Description
As shown in Table 3, we give the number of samples in the data set HC3-SI. For Chinese data, the number of samples in the train/val/test set are . For English data, the number of samples in the train/val/test set The sample sizes are respectively. The total sample size of our proposed HC3-SI is , nearly twice that of HC3. In addition, our data does not overlap with HC3, and they can be used together.